A woman invested 4 times as much at 8% as she did at 7%. The total amount of interest she earns in 1 year from both accounts is $2,730. How much did she invest at each rate?
Answer by mananth(16949)

let her invest $ x at 7%
She will invest $4x at 8%
...
Interest = 2,730
0.07x+0.08*4x=2730
0.07x+0.32x=2730
0.39x=2730
x= $7000 at 7%
So she would have invested 4*7000 = $28000 at 8%
...
m.ananth@hotmail.ca